# EgoBlind: Towards Egocentric Visual Assistance for the Blind People

### v1

@@ Abstract
We present EgoBlind, the first egocentric VideoQA dataset collected from blind individuals to evaluate the assistive capabilities of contemporary multimodal large language models (MLLMs). EgoBlind comprises 1,210 videos that record the daily lives of real blind users from a first-person perspective. It also features 4,927 questions directly posed or generated and verified by blind individuals to reflect their needs for visual assistance under various scenarios. We provide each question with an average of 3 reference answers to alleviate subjective evaluation. Using EgoBlind, we comprehensively evaluate 15 leading MLLMs and find that all models struggle, with the best performers achieving accuracy around 56%, far behind human performance of 87.4%. To guide future advancements, we identify and summarize major limitations of existing MLLMs in egocentric visual assistance for the blind and provide heuristic suggestions for improvement. With these efforts, we hope EgoBlind can serve as a valuable foundation for developing more effective AI assistants to enhance the independence of the blind individuals’ lives.
@@ 1 Introduction
The rapid advancement of multimodal large language models (MLLMs) [1, 20, 26, 35, 29, 34] has significantly improved the performance of visual question answering (VQA). However, existing VQA datasets primarily focus on third-person perspective [12, 37] or general-purpose image and video understanding [17, 6, 28], overlooking critical real-world applications such as visual assistance for the visually-impaired individuals who constitute an overwhelming population of 2.2 billion people around the world [30, 14]. The research is especially scarce in assisting the blind from a first-person perspective [33].
Figure 1: Real example of EgoBlind about blind user seeking egocentric visual assistance. The first time when the user puts their hand on the knob of the microwave, they are curious about what the knob control. The second time after they tune the knob, they are interested to know the position of the knob. The third time when they put their hand on another knob, they are curious about what knob it is. Finally, they want to know if it is correct to turn the knob clock-wise, to set the cooking time.
@@
In light of this, we construct EgoBlind – the first egocentric VideoQA dataset designed to benchmark and advance MLLMs towards live egocentric visual assistance for blind people. EgoBlind comprises 1,210 egocentric videos that capture the visual experiences of blind individuals and 4,927 questions that are directly posed or automatically generated and further verified by blind users to reflect their assistive needs in exploring their surroundings. To categorize these needs for better analysis, the questions are grouped into six key types: Information Reading, Safety Warning, Navigation, Social Communication, Tool Use, and Other Resources. We set the QA task as real-time (timestamp restricted) and open-form answer generation to better align with its live assistance nature. We provide multiple reference answers for each question for effective evaluation and carefully prompt the LLM for judgment.
@@
By focusing on blind users’ needs for real-time egocentric visual assistance, we aim to explore several key challenges in video-language learning, including egocentric dynamic scene understanding with poor visual quality (e.g., unstable motion, object blur, and occlusions), real-time context-aware user intention reasoning, and assistance-oriented answer generation. Take Fig. 1 as an example; engaging in the user’s activity from a first-person perspective is crucial to understanding and answering the questions. This specification requires capturing the gaze area and hand motion and reasoning the related visual content spatio-temporally. Finally, this should be achieved with poor quality visual inputs, as the regions of interest are often not in focus (e.g., the control panel in Fig. 1).
@@
With EgoBlind, we benchmark 15 recent MLLMs, covering both the open-source models (e.g., InternVL2 [5], Qwen2-VL [35] and LLaVA-OV [19] and VideoLLaMA3 [44]) and the closed-source ones (e.g., GPT-4o [29] and Gemini 2.0 [34]). For open-source models, we specially take into account the models that 1) achieve the state-of-the-art (SOTA) results on common video QA benchmarks [43, 37, 27, 11], and benchmarks emphasizing egocentric understanding [28, 6, 42]. Additionally, we include three SOTA modes (Video-LLaVA [22] and LLaMA-VID [21], VILA [23]) on the image blind QA dataset VizWiz [14].
@@
  • Models are poor in assisting with egocentric navigation, safety warning, and tool use, indicating limited knowledge in these fields.
  • Models generate objectively accurate answers about visual content, but the answers often fail to meet blind users’ subjective needs for assistance. This result suggests a weakness in reasoning about user intentions within dynamic egocentric visual contexts.
  • Models, especially the open-source ones, are sycophantic [47]; they hallucinate wrong and potentially malicious answers when the blind users ask about non-existent objects in their surroundings due to blindness.
  • Models fail to reason the temporal contexts and localize objects relative to the blind users’ real-time locations.
Our contributions are: 1) we tackle significant challenges to construct the first VideoQA dataset (EgoBlind) to benchmark and promote research toward egocentric visual assistance for the blind, and 2) we comprehensively analyze the behaviors of the leading MLLMs, revealing lots of limitations and possibilities for improvements.
@@ 2.1 VQA for the Blind
Visual question answering (VQA) is the task of answering user questions about images and videos. As one of its most impactful applications, VQA has been proposed to help visually impaired individuals [2]. Yet, so far the vast majority of advancements have been made in general-purpose VQA [12, 40, 37, 43] where the questions are to test the models’ capability of general-purpose visual understanding, with also the images and videos captured from third-person perspective. The closest in aim is VizWiz [14], which collects data from real human-powered VQA systems for helping blind users (e.g., BeMyEye [^1] http://www.bemyeyes.org/). Despite the significance, VizWiz limits in static images and object-centric questions; it fails to cater to the real-time and broader assistive needs of blind users in exploring the dynamic surroundings. Recent ongoing work VIEW-QA [33] proposes to capture the daily challenges faced by visually impaired individuals via using 360-degree egocentric wearable cameras. However, its videos and questions are collected by training actors to simulate the experiences of blind individuals, limiting its ability to reflect the real blind people’s daily lives and their true needs for visual assistance, especially for those who have been blind for a long time.
2.2 Egocentric VQA
@@
Egocentric VQA has earned increasing interest for its significant application value towards embodied assistance [31, 39]. Early effort EgoVQA [10] is small scale and limits its questions in challenging action recognition from first-person perspective with the absence of the camera wearer in the footage. Subsequent advancement EgoTaskQA [17] uses machine-generated questions to evaluate models’ task understanding capabilities. AssistQ [36] is close to our aim for embodied assistance, but it limits to instructional videos for demonstration of tool use. Ever since the release of Ego4D [13], lots of real-world egocentric QA tasks are explored [24, 32, 3, 28, 9, 7, 6, 42]. However, most of they aim at general-purpose egocentric visual understanding, or focus on a specific aspect of assistance, e.g., QAEgo4D [3] for episodic memory and EgoTextVQA [48] for scene text understanding. To our best knowledge, there is so far no existing egocentric VQA dataset specially collected from real blind individuals.
2.3 MLLMs for VQA
@@
MLLMs have significantly improved the landscape of VQA. By integrating visual information into powerful LLMs through lightweight connection modules (e.g., MLP [25]), MLLMs extend the capabilities of LLMs to converse with images and videos naturally like humans. This brings breakthrough over traditional VQA which answers questions by either multi-choice selection [16, 37] or close-vocabulary classification [40, 43]. While most of the MLLMs are developed for VQA of third-person perspective [39, 27, 45, 22, 38, 21, 8, 35, 44], recent advancements [24, 32, 46, 4, 5, 42] specially include ego videos for understanding. Nonetheless, all of they target a general-purpose egocentric visual understanding. In this paper, we will comprehensively examine their capacities towards egocentric visual assistance for the blind users.
@@ 3.1 Dataset Creation
Initially, we have tried to invite blind volunteers to record their daily lives from a first-person perspective and pose questions about their needs for daily visual assistance, similar to the data collection process of Ego4D [13]. However, we find that this approach presents significant challenges regarding data diversity (we can only get contact to a very limited number of blind people), quality (people often sparsely and coarsely ask what surrounds them to complete the data collection task), as well as ethical and privacy concerns. As an alternative, we collect videos from social media platforms and adopt the annotation pipeline as illustrated in Fig. 2. We elaborate on the details below.
@@ 3.1.1 Video Collection
We crawl 425 long-form egocentric videos from video media platforms such as Bilibili and TikTok, all originating from visually impaired content creators. These videos, captured from the first-person perspective of blind individuals using GoPro or mobile phone, document their daily lives (such as in traveling, cooking, navigating, social gatherings, medical visiting.) and provide real-world data on how visually impaired individuals perform tasks and solve problems in dynamic environments. Since a single long video often contains a vast array of visual content, e.g., a summarized record of a visually impaired creator’s week, we manually parition the videos into distinct segments following the traces of splicing. During editing, we delete video moments with sharp scene transitions and big subtitles to enhance the focus of each clip. Ultimately, we obtain 1,210 video clips with an average duration of 35.8 seconds.
@@ 3.1.2 Question Collection
We obtain the questions via three different approaches: 1) Manual Extraction: We manually extract the visual-assistance related questions directly posed by the blind users in the videos. 2) AI-Generation: We prompt GPT-4o [29] to act as blind user and generate questions by engaging in their egocentric perspective. The generated questions are further verified and edited as necessary by both sighted and visually impaired people, ensuring their correctness and alignment with blind users’ needs. 3) Blind User Annotation: For some videos, we introduce their major contents to the blind individuals and invite them to annotate questions when confronted with the visual situations. The actual annotation is conducted within the blind user verification stage. The collection details are as follows:
@@
Information Reading 2283 / 962 Q: What floor is the elevator currently on?
@@
1181 / 629 Q: Is it safe to cross the street now?
@@
Navigation 669 / 389 Q: Where is the entrance to the building?
@@
366 / 258 Q: Is there anyone nearby that I can ask for directions?
@@
Manual Extraction. We first manually watch the video clips and extract the vision-assistive questions from the camera wearers, the timestamps of the questions are also recorded. It is worth noting that these in-video questions are answered by the blind users’ sighted partners or others in the videos. Thus, to prevent answer leakage and ensure real-time QA, models can only access visual content up to the question timestamps to answer a specific question. We obtain a total of 500 questions during this process since qualified questions are sparse in the videos.
AI-Generation. To enrich the annotations, we consider generating questions by prompting GPT-4o [29] to act as blind questioners, followed by rigorous verification with real blind individuals. Specifically, we first categorize the assistance scenarios into six types based on our observation of the video content and the extracted questions: Information Reading, Safety Warning, Navigation, Social Communication, Tool Use, and Other Resources. Typical question of each group is listed in Tab. 1. To limit generative hallucinations and ensure a diverse set of examples for each assistance type, QA generation is conducted categorically, focusing on one assistance type at a time. Moreover, the generation and verification are conducted alternately to ensure high generation quality.
Specifically, in our prompting, the generated QA pairs are required to adhere to the following criteria to ensure that they are practical, context-aware, and closely aligned with the needs of the visually impaired individuals: 1) First-Person Perspective: All questions should be framed from the perspective of the visually impaired individuals (camera wearers). 2) Practicality: The questions must be practical and reflect the blind’s true needs in the given situation. 3) Video-Level Questions: Questions should reflect the temporal aspect of the video, requiring multiple frames to derive the answer and emphasizing dynamic characteristics. 4) Real-Time Context: Questions must pertain to the real-time context of the video and exclude any external information beyond the visible content before and at the question time stamp. Other details are presented in Appendix A.
@@ 3.1.3 Manual Verification
Our verification is done in three stages involving both sighted and visually impaired people. Also, the first three authors are engaged into all stages to control the quality. Concretely, in the first stage, the authors are responsible to check the QA quality every after generating a limited number of questions for one type of assistance. Related issues are recorded and reflected in the prompts for subsequent generations, thus to maximally ensure the generation quality. In such an alternate way, we generate 16,217 questions.
@@
In the second stage, we invite 63 volunteers (sighted people) to thoroughly review the questions. In this process, we remove lots of questions that are 1) redundant in meaning; 2) not easy to answer (e.g., “How do I use my phone?”); and 3) meaningless (e.g., “How do I use my white cane?”). We also find and delete a moderate number of questions that are 1) not related to vision (e.g., “How heavy is this box?”) and 2) not from first-person perspective (e.g., “what is the man (camera wearer) doing?”) . Note that all ambiguous questions are further cross-validated by the three authors to either remove or edit them. What remained are 4,737 questions after this stage, in which 82.8% are edited by humans. Additionally, the 500 extracted questions also undergo checking for translation errors, resulting in a total of 5,237 valid questions after this stage.
@@
In the third stage, we sample 30% data (1.5K) corrected questions and invite 111 visually impaired volunteers (with diverse backgrounds) to review them to ensure their relevance to the blind users’ true needs for visual assistance (see Appendix A). The result observations are then reflected throughout the whole dataset by the authors. Specifically, the blind volunteers are requested to score the confidence (from the lowest 0 to the highest 5) of a question that it may be asked under a specific visual scenario. Each question will be scored by at least 10 volunteers and maximally 33. Additionally, the bind volunteers are encouraged to pose new questions optionally. With the study result, we further remove the questions with an average confidence score of lower than 1 and add a limited number of new questions (about 300) annotated by the blind volunteers. Other details are presented in Appendix A. Eventually, we obtain 4,927 valid questions. Their distributions are listed in Tab. 1.
3.1.4 Answer Annotation

### v2

@@ Abstract
We present EgoBlind, the first egocentric VideoQA dataset collected from blind individuals to evaluate the assistive capabilities of contemporary multimodal large language models (MLLMs). EgoBlind comprises 1,392 videos that record the daily lives of real blind users from a first-person perspective. It also features 5,311 questions directly posed or generated and verified by blind individuals to reflect their in-situation needs for visual assistance under various scenarios. We provide each question with an average of 3 reference answers to alleviate subjective evaluation. Using EgoBlind, we comprehensively evaluate 16 advanced MLLMs and find that all models struggle, with the best performers achieving accuracy near 60%, far behind human performance of 87.4%. To guide future advancements, we identify and summarize major limitations of existing MLLMs in egocentric visual assistance for the blind and explore heuristic solutions for improvement. With these efforts, we hope EgoBlind can serve as a valuable foundation for developing more effective AI assistants to enhance the independence of the blind individuals’ lives. Data and evaluation code are available at https://github.com/doc-doc/EgoBlind.
@@ 1 Introduction
The rapid advancement of multimodal large language models (MLLMs) [1, 2, 3, 4, 5, 6] has significantly improved the performance of visual question answering (VQA). However, existing VQA datasets primarily focus on third-person perspective [7, 8] or general-purpose image and video understanding [9, 10, 11], overlooking critical real-world applications such as visual assistance for the visually-impaired who constitute an overwhelming population of 2.2 billion people around the world [12]. The research is especially scarce in assisting the blind from a first-person perspective [13].
@@
In light of this, we construct EgoBlind – the first egocentric VideoQA dataset designed to benchmark and advance MLLMs towards live egocentric visual assistance for the blind. EgoBlind comprises 1,392 egocentric videos that capture the visual experiences of blind individuals and 5,311 questions that are directly posed or automatically generated and further verified by blind users to reflect their assistive needs in exploring their surroundings. To categorize these needs for better analysis, the questions are grouped into six key types: “Information Reading”, “Safety Warning”, “Navigation”,“Social Communication”, “Tool Use”, and “Other Resources”. We set the QA task as real-time (timestamp restricted) and open-form answer generation to better align with its live assistance nature. Multiple reference answers with well-aligned evaluation prompts are also provided for effective assessment.
Figure 1: Real example of EgoBlind about a blind user seeking egocentric visual assistance. The first time when she puts her hand on the knob of the microwave, she is curious about what the knob control. The second time after she tunes the knob, she is interested to know the position of the knob. The third time when she puts her hand on another knob, she is curious about what knob it is. Finally, she wants to know if it is correct to turn the knob clock-wise, to set the cooking time.
@@
By focusing on blind users’ needs for real-time egocentric visual assistance, we aim to explore several key challenges in video-language learning, including egocentric dynamic scene understanding with poor visual quality (e.g., unstable motion, object blur, and occlusions), in-situation user intention reasoning, assistive and reliable answer generation. Take Figure 1 as an example; engaging in the user’s activity from a first-person perspective is crucial to understanding and answering the questions. This specification requires capturing the gaze area and hand motion and reasoning the related visual content spatio-temporally. Finally, this should be achieved with poor quality visual inputs, as the regions of interest are often not in focus (e.g., the control panel in Figure 1).
@@
With EgoBlind, we benchmark 16 recent MLLMs, covering both the open-source models (e.g., InternVL2.5 [14], Qwen2.5-VL [4], LLaVA-OV [15] and VideoLLaMA3 [16]) and the closed-source ones (e.g., GPT-4o [5] and Gemini [6]). For open-source models, we specially take into account the models that 1) achieve the state-of-the-art (SOTA) results on common video QA benchmarks [17, 8, 18, 19], and benchmarks emphasizing egocentric understanding [11, 10, 20]. Additionally, we include three modes (Video-LLaVA [21] and LLaMA-VID [22], VILA1.5 [23]) that achieve SOTA on the image blind QA dataset VizWiz [24].
@@
  • Models are poor in assisting with egocentric navigation, safety warning, and tool use (open-source ones), indicating limited knowledge in these fields.
  • Models can generate objectively accurate answers relative to the visual contents, but the answers often fail to meet blind users’ subjective needs for assistance, indicating a weakness in reasoning about user intentions in specific situations.
  • Models are struggling in understanding the spatio-temporal contexts and localizing objects relative to the users’ real-time locations.
  • Models, especially the open-source ones, are sycophantic [25]; they hallucinate wrong and potentially malicious answers when the users’ questions deviate the visual facts due to blindness.
To summarize our major contributions, 1) we tackle significant challenges to construct the first VideoQA dataset (EgoBlind) to benchmark and promote research toward egocentric visual assistance for the blind, and 2) we comprehensively analyze the behaviors of the leading MLLMs, revealing lots of limitations and possibilities for improvements.
@@ 2 Related Work
VQA for the Blind. Visual question answering (VQA) is the task of answering user questions about images and videos. As one of its most impactful applications, VQA has been proposed to help visually impaired individuals [26]. Yet, so far the vast majority of advancements have been made in general-purpose VQA [7, 27, 8, 17] where the questions are to test the models’ capability of general-purpose visual understanding, with also the images and videos captured from third-person perspective. The closest in aim is VizWiz [24], which collects data from real human-powered VQA systems for helping blind users (e.g., BeMyEye [^1] http://www.bemyeyes.org/). Despite the significance, VizWiz limits in static images and object-centric questions; it fails to cater to the real-time and broader assistive needs of blind users in exploring the dynamic surroundings. Recent ongoing work VIEW-QA [13] proposes to capture the daily challenges faced by visually impaired individuals via using 360-degree egocentric wearable cameras. However, its videos and questions are collected by training actors to simulate the experiences of blind individuals, limiting its ability to reflect the real blind people’s daily lives and their true needs for visual assistance, especially for those who have been blind for a long time.
@@
Egocentric VQA. Egocentric VQA has earned increasing interest for its significant application value towards embodied assistance [28, 29]. Early effort EgoVQA [30] is small scale and limits its questions in challenging action recognition from first-person perspective with the absence of the camera wearer in the footage. Subsequent advancement EgoTaskQA [9] uses machine-generated questions to evaluate models’ task understanding capabilities. AssistQ [31] is close to our aim for embodied assistance, but it limits to instructional videos for demonstration of tool use. Ever since the release of Ego4D [32], lots of real-world egocentric QA tasks are explored [33, 34, 35, 11, 36, 37, 10, 20]. However, most of they aim at general-purpose egocentric visual understanding, or focus on a specific aspect of assistance, e.g., QAEgo4D [35] for episodic memory, EgoTextVQA [38] for scene text understanding and EgoLifeQA [39] for long context life assistants. To our best knowledge, there is so far no existing egocentric VQA dataset specially collected from real blind individuals.
@@
MLLMs for VQA. MLLMs have significantly improved the landscape of VQA. By integrating visual information into powerful LLMs through lightweight connection modules (e.g., MLP [40]), MLLMs extend the capabilities of LLMs to converse with images and videos naturally like humans. This brings breakthrough over traditional VQA which answers questions by either multi-choice selection [41, 8] or close-vocabulary classification [27, 17]. While most of the MLLMs are developed for VQA of third-person perspective [29, 18, 42, 21, 43, 22, 44, 4, 16], recent advancements [33, 34, 45, 46, 14, 20] specially include ego videos for understanding. Nonetheless, all of they target a general-purpose egocentric visual understanding. In this paper, we will comprehensively examine their capacities towards egocentric visual assistance for the blind users.
@@ 3.1 Dataset Creation
Initially, we have tried to invite blind volunteers to record their daily lives from a first-person perspective and pose questions about their needs for daily visual assistance. However, we find that this approach presents significant challenges regarding data diversity (we can only get contact to a very limited number of blind people), quality (people often sparsely and coarsely ask what surrounds them to complete the data collection task), as well as ethical and privacy concerns. As an alternative, we collect videos from video media platforms and adopt the annotation pipeline as illustrated in Figure 2. We elaborate on the details below.
@@ 3.1.1 Video Collection
We search and crawl 478 long-form egocentric videos (in batches) from video media platforms such as Bilibili and TikTok, ensuring that all videos are originated from visually impaired content creators. These videos, captured from the first-person perspective of blind individuals using GoPro or mobile phone, document their daily lives (such as in traveling, cooking, navigating, shopping, social gatherings, medical visiting.) and provide real-world data on how visually impaired individuals perform tasks and solve problems in dynamic environments. Since a single long video often contains a vast array of visual content, e.g., a summarized record of a visually impaired creator’s week, we manually parition the videos into distinct segments following the traces of splicing. During editing, we delete video moments with sharp scene transitions and big subtitles to enhance the focus of each clip. After editing, we obtain 1,329 video clips with an average duration of 40.0 seconds.
@@ 3.1.2 Question Collection
We obtain the questions via three different approaches: 1) Manual Extraction: We manually extract the visually-assistive questions directly posed by the blind users in the videos. 2) AI-Generation: We prompt GPT-4o [5] to act as blind users and generate questions by engaging in their egocentric perspective. The generated questions are further verified and edited by both sighted and visually impaired people, ensuring their correctness and alignment with blind users’ true needs. 3) Blind User Annotation: For some videos, we introduce their major contents to the blind individuals and invite them to annotate assistive questions when confronted with the visual situations. The actual annotation is done together with the verification stage. Our collection details are as follows:
Manual Extraction. We first manually watch the video clips and extract the vision-assistive questions from the camera wearers, the timestamps of the questions are also recorded. It is worth noting that these in-video questions are answered by the blind users’ sighted partners or others in the videos. Thus, to prevent answer leakage and ensure real-time QA, models can only access visual content up to the question timestamps to answer a specific question. We obtain a total of 541 questions during this process since qualified questions are sparse in the videos.
AI-Generation. To enrich the annotations, we consider generating questions by prompting GPT-4o [5] to act as blind questioners, followed by rigorous verification with real blind individuals. Specifically, we first categorize the assistance scenarios into six types based on our observation of the video content and the extracted questions: Information Reading, Safety Warning, Navigation, Social Communication, Tool Use, and Other Resources. Typical question of each group is listed in Table 1. To ensure a diverse set of examples for each assistance type, the generation is conducted categorically with tailored prompt for each question type. In addition, we also prompt GPT to generate a reference answer for each question.
Specifically, in our prompting for each category, the generated QA pairs are required to adhere to the following criteria to ensure that they are practical, context-aware, and closely aligned with the needs of the blind: 1) First-Person Perspective: All questions should be framed from the perspective of the visually impaired individuals (camera wearers). 2) Practicality: The questions must be practical and reflect the blind’s true needs in the given situation. 3) Video-Level Questions: Questions should reflect the temporal aspect of the video, requiring multiple frames to derive the answer and emphasizing dynamic characteristics. 4) Real-Time Context: Questions must pertain to the real-time context of the video and exclude any external information beyond the visible content up to the question time stamp. Other details are presented in Appendix A.1.
@@
Information Reading 2,464 / 1,040 Q: What floor is the elevator currently on?
@@
1,260 / 686 Q: Is it safe to cross the street now?
@@
Navigation 751 / 438 Q: Where is the entrance to the building?
@@
395 / 280 Q: Is there anyone nearby that I can ask for directions?
@@ 3.1.3 Manual Verification
Our verification is done in three stages involving both sighted and visually impaired people. The first three authors are engaged into all stages to control the quality. Concretely, in the first stage, the authors are responsible to check the QA quality every after generating a limited number of questions for one type of assistance. Related issues are recorded and reflected in the prompts for subsequent generations, thus to maximally ensure the generation quality. In such an alternate way, we generate 16,560 questions in batches.
@@
In the second stage, we invite 63 volunteers (sighted people) to thoroughly review the questions. In this process, we remove lots of questions that are 1) redundant in meaning; 2) not easy to answer (e.g., “How do I use my phone?”); and 3) meaningless (e.g., “How do I use my white cane?”). We also find and delete a moderate number of questions that are 1) not related to vision (e.g., “How heavy is this box?”) and 2) not from first-person perspective (e.g., “what is the man (camera wearer) doing?”) . Note that all ambiguous questions are further cross-validated by the three authors to either remove or edit them. What remained are 5,080 questions after this stage, in which more than 80% of data are edited by humans. Additionally, the 541 extracted questions also undergo checking for translation errors, resulting in a total of 5,621 valid questions after this stage.
@@
In the third stage, to reduce the annotation burden for the blind, we sample 30% (1.5K) questions and invite a total of 111 visually impaired volunteers to review them to ensure their relevance to the blind users’ true needs for visual assistance (details are presented in Appendix A.2). The result observations are then reflected throughout the whole dataset by the authors. Specifically, the blind volunteers are requested to score the confidence (from the lowest 0 to the highest 5) of a question that it may be asked under a specific visual scenario. Each question will be scored by at least 10 volunteers and maximally 33. Additionally, the bind volunteers are encouraged to pose new questions optionally. With the study result, we further remove the questions with an average confidence score of lower than 1 and add a limited number of new questions (about 300) annotated by the blind volunteers. Eventually, we obtain 5,311 valid questions. Their distributions are listed in Table 1.
3.1.4 Multiple Answer Annotation